# Probing Primordial Black Hole Formation from Domain Wall Isocurvature Perturbations: Constraints and Implications

### v2

@@ 3 Density fluctuations from domain walls
After the production of domain walls via the Kibble mechanism [6], the size of the walls rapidly grows to a superhorizon scale through the merger and expansion of the walls. Subsequently, the walls re-enter the horizon due to the combined effects of chopping and the pressure from the bias potential. During the horizon crossing of the domain walls, the network ultimately becomes a few-body system composed of individual domain walls. Consider a domain wall network consisting of $N$ identical walls, each with mass $M_{w}$, randomly distributed within the Hubble volume $V_{H}\equiv 4\pi r_{H}^{3}/3$, where $r_{H}=1/H$. We employ the Poisson distribution to describe the probability of finding $n$ walls within a volume of $V\equiv 4\pi r^{3}/3$
@@
where $V_{w}=4\pi\bar{L}^{3}/3$ is the volume of a single wall, and $V=nV_{w}$. The applicability of the Poisson distribution to the domain wall network has been validated through simulations in Refs. [51, 52]. The spatially averaged energy density of the domain wall within the volume $V$ is given by $\bar{\rho}_{w}(r)=nM_{w}/V$. The ensemble average of $\bar{\rho}_{w}(r)$ is then expressed as
@@
$\langle(\delta\bar{\rho}_{w}(r))^{2}\rangle=\langle\rho_{w}^{2}(r)\rangle-\langle\bar{\rho}_{w}\rangle^{2}=\sum_{n=0}^{\infty}\rho_{w}^{2}(r)\mathbb{P}_{n}(V)-\langle\bar{\rho}_{w}\rangle^{2}=\frac{M_{w}^{2}}{V_{w}V}.$ (3.4)
@@
where we have used the relations $f_{w}(t)=\langle\bar{\rho}_{w}\rangle/\rho_{\rm rad}\simeq\langle\bar{\rho}_{w}\rangle/\rho_{\rm tot}$ and $V_{H}=NV_{w}$.
@@
$\bar{\rho}_{w}(r)=\frac{1}{V}\int_{V}\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}\rho_{w}(\bm{x})\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ \overline{\delta\rho}_{w}(r)=\frac{1}{V}\int_{V}\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}\delta\rho_{w}(\bm{x}).$ (3.7)
@@
$\left\langle\int_{V}\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}\int_{V}\mathrm{\leavevmode\nobreak\ d}^{3}\bm{x}^{\prime}\delta\rho_{w}(\bm{x})\delta\rho_{w}(\bm{x}^{\prime})\right\rangle=\frac{M_{w}^{2}V}{V_{w}}.$ (3.8)
The randomness of the spatial distribution implies that fluctuations at two different locations are independent [53]. Furthermore, the spatial distribution is assumed to be homogeneous and isotropic. With these properties, we can express the density correlator as
@@
Simulations [54, 55, 49] indicate that the domain walls tend to merge with the neighboring ones in the evolution, resulting in a wall network with a small number of objects. For a model with a $Z_{2}$ symmetry, there eventually exists $N=1$ domain wall that separates two domains [54]. For models with $\mathcal{N}\sim 10$ domains, one expects that $N\sim\mathcal{N}$ domain walls exist in the late stage of the network evolution. Therefore, once the domain walls contribute significantly to the total energy density, with $f_{w}\sim 0.1$, near the time of domain wall annihilation, the energy density perturbation arising from the number density fluctuations of the domain walls can become remarkable.
@@ 4 Isocurvature perturbation from domain wall network
where the sound speed $c_{s}^{2}=\bar{p}^{\prime}/\bar{\rho}^{\prime}$. In Fourier space, the Einstein equations for scalar perturbations are given by [56]
@@
$\left(\frac{k}{\mathcal{H}}\right)^{2}(\Psi-\Phi)=3w\Pi,$
$\mathcal{H}^{-1}\Psi^{\prime}+\Phi=\frac{3}{2}(1+w)\frac{\mathcal{H}}{k}v^{N},$
@@
$\omega=\frac{\sum\bar{\rho}_{i}\omega_{i}}{\bar{\rho}},\leavevmode\nobreak\ \leavevmode\nobreak\ c_{s}^{2}=\sum\frac{\bar{\rho}_{i}+\bar{p}_{i}}{\bar{\rho}+\bar{p}}c_{i}^{2},\leavevmode\nobreak\ \leavevmode\nobreak\ \delta=\frac{\sum\bar{\rho}_{i}\delta_{i}}{\bar{\rho}},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ \leavevmode\nobreak\ v=\sum\frac{1+w_{i}}{1+w}\frac{\bar{\rho}_{i}}{\bar{\rho}}v_{i}.$ (4.12)
@@
$S^{\prime\prime}=-kv_{\rm rel}^{\prime}=$ $\leavevmode\nobreak\ k\mathcal{H}\left[\left(1-3w_{i}\right)v_{i}^{N}-\left(1-3w_{j}\right)v_{j}^{N}\right]+k\left(\frac{w_{i}^{\prime}}{1+w_{i}}v_{i}^{N}-\frac{w_{j}^{\prime}}{1+w_{j}}v_{j}^{N}\right)$ (4.16)
@@
$\left\{\begin{matrix}a_{11}=\frac{(1+\omega_{i})(1+f)}{\lambda},\leavevmode\nobreak\ \\ a_{12}=\frac{(1+\omega_{i})(1+\omega_{j})}{\lambda},\end{matrix}\right.\quad\quad{\rm and}\quad\quad\left\{\begin{matrix}a_{21}=-\frac{(1+\omega_{j})(1+f)}{\lambda},\\ a_{22}=\frac{(1+\omega_{i})(1+\omega_{j})f}{\lambda},\end{matrix}\right$. (4.18)
@@
$\left\{\begin{matrix}b_{11}=\frac{(1+\omega)(1+f)}{\lambda},\\ b_{12}=\frac{1+\omega_{j}}{\lambda},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \end{matrix}\right.\quad\quad{\rm and}\quad\quad\left\{\begin{matrix}b_{21}=\frac{(1+\omega)(1+f)}{\lambda},\\ b_{22}=\frac{(1+\omega_{i})f}{\lambda},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \end{matrix}\right$. (4.19)
@@
If the components are cold matter and radiation, the parameters reduce to $\mu_{1}=-1/(1+w)$ and $\mu_{2}=(1-3c_{s}^{2})/3$. In this case, we reproduce the so-called Kodama-Sasaki equation [57].
@@ 5 Critical collapse
Let us assume that the perturbed region is spherically symmetric and has an initial radius $R(t)$ larger than the Hubble scale. The locally overdense region can then constitute a spatially closed Universe with a metric given by [58, 59]
@@
The initial expansion rates of the background and the perturbed region can be set to be the same by choosing the coordinates [60, 59] so that the initial density contrast is given by
@@
Together with Eqs. (5.5) and (5.6), we obtain the critical scale factor that the overdense region stops expanding
@@
The overdense region could contrast further against the pressure gradient if the self-gravitation $\Phi$ of the mass within the critical radius $R_{c}$ is larger than the internal energy $U$ of the overdense region, which gives [33]
@@ 6.1 PBH relic abundance
It is expected that the PBHs formed from the Hubble horizon (re-)entry of the large density perturbations could have masses at the same order as the associated horizon mass $M_{H}\simeq 2\times 10^{15}(t/1s)M_{\odot}$. To further estimate the PBH production, let’s take the naive relation between the PBH mass $M$ and the horizon mass from critical phenomena [37, 38]
@@
Taking advantage of the Press-Schechter formalism, the PBH energy fraction can be expressed as [61]
@@
where we consider a Gaussian probability distribution (3.19) for the density contrast. Note that we use an upper limit of infinity for the integration and confirm that it is not different from the integration result with an upper limit of $\delta=1$ from the requirement of avoiding separate Universes [33, 59]. A recent work [62] argued that this upper limit was a result of the geometry of the selected slicing rather than as a consequence of avoiding separate Universes, which instead led to a divergence of density fluctuations on the superhorizon scale. With the PBH mass given by Eq. (6.2), the fraction of PBHs at the formation time is
@@
where ${\rm erfc}(x)$ is the complementary error function (see appendix D for more details). The critical density contrast $\delta_{c}$ varies with the equation of state $\omega$. For radiation phase $\delta_{c}=0.453$. We determine $\delta_{c}$ with the $\delta_{c}-\omega$ relation revealed by the simulations in Ref. [63] (see section 7).
A more precise PBH mass formation was obtained from the numerical computations on the collapse of a relativistic fluid, which indicated a simple scaling law for the PBH mass distribution function near the critical threshold [64, 65, 66]
@@
where the constant parameter $C=3.3$ and $b=0.36$ for the formation of PBH during radiation domination [67]. Note that these parameters may be affected by the phase of cosmic matter. Then the PBH energy fraction is given by
@@
$s_{f}a_{f}^{3}=s_{0}a_{0}^{3},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ \leavevmode\nobreak\ s(T)=\frac{2\pi^{2}}{45}g_{*,s}T^{3},$ (6.9)
where the current Universe entropy density $s_{0}=2891.2\leavevmode\nobreak\ \rm cm^{-3}$. Together with Eqs. (6.1)-(6.3) we obtain the the current PBH relic density abundance
@@
where the Hubble expansion rate is $H_{0}=100h\leavevmode\nobreak\ \mathrm{km\leavevmode\nobreak\ s}^{-1}\leavevmode\nobreak\ \mathrm{Mpc}^{-1}$ with $h=0.68$ from the Planck observation [68]. With the relation $M=\gamma_{p}M_{H}$ and the approximation $g_{*}(T)\simeq g_{*,s}(T)$, we reproduce $\Omega_{\rm PBH}$ given in Ref. [69]. We observe that the PBH energy fraction at the formation should be of order $\beta(M)\sim 10^{-9}$ so that the PBH relic abundance today does not exceed the required dark matter relic abundance. Thus the variance of the density contrast should be small and the probability density distribution should be concentrated so that the probability for the production of large density contrasts is exponentially suppressed in the tail of the probability density distribution.
@@ 6.2 Density fluctuation from inflation-induced curvature perturbation
Let’s first briefly review the PBH production in the inflation scenario. For the density fluctuations from the curvature perturbations during inflation, the variance is in general estimated by Eq. (3.17). On comoving hypersurfaces, the density perturbation can be related to the curvature perturbation by [70]
@@
where the amplitude $A_{\mathcal{R}_{c}}$ is a constant that depends on the inflation physics, and $k_{*}$ is the reference scale. Current CMB and LSS observations on large scales ($k\lesssim{\rm Mpc^{-1}}$) have restricted the primordial curvature power spectrum at the order around $10^{-9}$. To have a sizable PBH formation for LIGO observation, the primordial curvature power spectrum should be of order $\mathcal{P}_{\mathcal{R}_{c}}(k)\sim 10^{-2}-10^{-1}$ (corresponding to a variance in the range $10^{-4}\lesssim\sigma^{2}\lesssim 10^{-2}$ which gives $\beta\sim 10^{-9}-10^{-6}$, see Fig. 11 in appendix D), which requires a blue-tilted spectral index $n_{s}\simeq 1.85$ at scale $k\sim 10^{6}\leavevmode\nobreak\ \rm Mpc^{-1}$ (corresponding to a PBH mass $M\sim 30M_{\odot}$[37, 67]. On the other hand, the CMB observations have constrained the spectral index to be red-tilted, $n_{s}=0.968\pm 0.006$ at the scale $k_{*}=0.05\leavevmode\nobreak\ \rm Mpc^{-1}$. The relatively large blue-tilted spectral index at a small scale cannot be realized in the single-field slow-roll inflation but may be achieved in the running mass inflation or multi-scalar inflation models, in which the scale-dependent spectral index can be realized (see Refs. [71, 72, 73, 74, 35, 75, 76] for example).

### v3

@@ 3 Density fluctuations from domain walls
After the production of domain walls via the Kibble mechanism [Kibble:1976sj], the size of the walls rapidly grows to a superhorizon scale through the merger and expansion of the walls. Subsequently, the walls re-enter the horizon due to the combined effects of chopping, Hubble friction, and the pressure from the bias potential. During the horizon crossing of the domain walls, the network ultimately becomes a few-body system composed of individual domain walls. Consider a domain wall network consisting of $N$ identical walls, each with mass $M_{w}$, randomly distributed within the Hubble volume $V_{H}\equiv 4\pi r_{H}^{3}/3$, where $r_{H}=1/H$. We employ the Poisson distribution to describe the probability of finding $n$ walls within a volume of $V\equiv 4\pi r^{3}/3$
@@
where $V_{w}=4\pi\bar{L}^{3}/3$ is the volume of a single wall, and $V=nV_{w}$. The applicability of the Poisson distribution to the domain wall network has been validated through simulations in Refs. [Takahashi:2020tqv, Kitajima:2022jzz]. The spatially averaged energy density of the domain wall within the volume $V$ is given by $\bar{\rho}_{w}(r)=nM_{w}/V$. The ensemble average of $\bar{\rho}_{w}(r)$ is then expressed as
@@
$\langle(\delta\bar{\rho}_{w}(r))^{2}\rangle=\langle\bar{\rho}_{w}^{2}(r)\rangle-\langle\bar{\rho}_{w}\rangle^{2}=\sum_{n=0}^{\infty}\bar{\rho}_{w}^{2}(r)\mathbb{P}_{n}(V)-\langle\bar{\rho}_{w}\rangle^{2}=\frac{M_{w}^{2}}{V_{w}V}.$ (3.4)
@@
where we have used the relations $f_{w}(t)=\langle\bar{\rho}_{w}\rangle/\rho_{\rm rad}\simeq\langle\bar{\rho}_{w}\rangle/\rho_{\rm tot}$ and $V_{H}=NV_{w}$. Our rigorous result in Eq. (3.6) confirms the density fluctuation $\delta\rho/\rho\sim G\sigma_{w}t$ estimated by Vilenkin (see Eq. (13.5.13) in Ref. [Vilenkin:2000jqa]) for a $Z_{2}$ network ($N\sim 1$), and thereby generalizes it to the case of an arbitrary domain wall number $N$.
@@
$\bar{\rho}_{w}(r)=\frac{1}{V}\int_{V}\mathrm{\penalty 10000\ d}^{3}\bm{x}\rho_{w}(\bm{x})\penalty 10000\ \penalty 10000\ {\rm and}\penalty 10000\ \penalty 10000\ \overline{\delta\rho}_{w}(r)=\frac{1}{V}\int_{V}\mathrm{\penalty 10000\ d}^{3}\bm{x}\delta\rho_{w}(\bm{x}).$ (3.7)
@@
$\left\langle\int_{V}\mathrm{\penalty 10000\ d}^{3}\bm{x}\int_{V}\mathrm{\penalty 10000\ d}^{3}\bm{x}^{\prime}\delta\rho_{w}(\bm{x})\delta\rho_{w}(\bm{x}^{\prime})\right\rangle=\frac{M_{w}^{2}V}{V_{w}}.$ (3.8)
The randomness of the spatial distribution implies that fluctuations at two different locations are independent [Papanikolaou:2020qtd]. Furthermore, the spatial distribution is assumed to be homogeneous and isotropic. With these properties, we can express the density correlator as
@@
Simulations [Press:1989yh, Hiramatsu:2012sc, Lazanu:2015fua] indicate that the domain walls tend to merge with the neighboring ones in the evolution, resulting in a wall network with a small number of objects. For a model with a $Z_{2}$ symmetry, there eventually exists $N=1$ domain wall that separates two domains [Press:1989yh]. For models with $\mathcal{N}\sim 10$ domains, one expects that $N\sim\mathcal{N}$ domain walls exist in the late stage of the network evolution. Therefore, once the domain walls contribute significantly to the total energy density, with $f_{w}\sim 0.1$, near the time of domain wall annihilation, the energy density perturbation arising from the number density fluctuations of the domain walls can become remarkable.
@@ 4 Isocurvature perturbation from domain wall network
where the sound speed $c_{s}^{2}=\bar{p}^{\prime}/\bar{\rho}^{\prime}$. In Fourier space, the Einstein equations for scalar perturbations are given by [Suonio:2014cp]
@@
$\left(\frac{k}{\mathcal{H}}\right)^{2}(\Psi-\Phi)=3w\Pi,$ (4.7)
$\mathcal{H}^{-1}\Psi^{\prime}+\Phi=\frac{3}{2}(1+w)\frac{\mathcal{H}}{k}v^{N},$ (4.8)
@@
$\omega=\frac{\sum\bar{\rho}_{i}\omega_{i}}{\bar{\rho}},\penalty 10000\ \penalty 10000\ c_{s}^{2}=\sum\frac{\bar{\rho}_{i}+\bar{p}_{i}}{\bar{\rho}+\bar{p}}c_{i}^{2},\penalty 10000\ \penalty 10000\ \delta=\frac{\sum\bar{\rho}_{i}\delta_{i}}{\bar{\rho}},\penalty 10000\ \penalty 10000\ {\rm and}\penalty 10000\ \penalty 10000\ v=\sum\frac{1+w_{i}}{1+w}\frac{\bar{\rho}_{i}}{\bar{\rho}}v_{i}.$ (4.12)
@@
$S^{\prime\prime}=-kv_{\rm rel}^{\prime}=$ $\penalty 10000\ k\mathcal{H}\left[\left(1-3w_{i}\right)v_{i}^{N}-\left(1-3w_{j}\right)v_{j}^{N}\right]+k\left(\frac{w_{i}^{\prime}}{1+w_{i}}v_{i}^{N}-\frac{w_{j}^{\prime}}{1+w_{j}}v_{j}^{N}\right)$ (4.16)
@@
$\left\{\begin{matrix}a_{11}=\frac{(1+\omega_{i})(1+f)}{\lambda},\penalty 10000\ \\ a_{12}=\frac{(1+\omega_{i})(1+\omega_{j})}{\lambda},\end{matrix}\right.\quad\quad{\rm and}\quad\quad\left\{\begin{matrix}a_{21}=-\frac{(1+\omega_{j})(1+f)}{\lambda},\\ a_{22}=\frac{(1+\omega_{i})(1+\omega_{j})f}{\lambda},\end{matrix}\right$. (4.18)
@@
$\left\{\begin{matrix}b_{11}=\frac{(1+\omega)(1+f)}{\lambda},\\ b_{12}=\frac{1+\omega_{j}}{\lambda},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \end{matrix}\right.\quad\quad{\rm and}\quad\quad\left\{\begin{matrix}b_{21}=\frac{(1+\omega)(1+f)}{\lambda},\\ b_{22}=\frac{(1+\omega_{i})f}{\lambda},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \end{matrix}\right$. (4.19)
@@
If the components are cold matter and radiation, the parameters reduce to $\mu_{1}=-1/(1+w)$ and $\mu_{2}=(1-3c_{s}^{2})/3$. In this case, we reproduce the so-called Kodama-Sasaki equation [Kodama:1986ud].
@@ 5 Critical collapse
Let us assume that the perturbed region is spherically symmetric and has an initial radius $R(t)$ larger than the Hubble scale. The locally overdense region can then constitute a spatially closed Universe with a metric given by [Carr:1975qj, Yokoyama:1995ex]
@@
The initial expansion rates of the background and the perturbed region can be set to be the same by choosing the coordinates [Harrison:1969fb, Yokoyama:1995ex] so that the initial density contrast is given by
@@
Together with Eqs. (5.5) and (5.6), we obtain the critical scale factor for which the overdense region stops expanding
@@
The overdense region could contrast further against the pressure gradient if the self-gravitation $\Phi$ of the mass within the critical radius $R_{c}$ is larger than the internal energy $U$ of the overdense region, which gives [Carr:1974nx]
@@ 6.1 PBH relic abundance
It is expected that the PBHs formed from the Hubble horizon (re-)entry of the large density perturbations could have masses at the same order as the associated horizon mass $M_{H}\simeq 2\times 10^{15}(t/1s)M_{\odot}$. To further estimate the PBH production, let’s take the naive relation between the PBH mass $M$ and the horizon mass from critical phenomena [Sasaki:2018dmp, Carr:2020gox]
@@
Taking advantage of the Press-Schechter formalism, the PBH energy fraction can be expressed as [Niemeyer:1997mt]
@@
where we consider a Gaussian probability distribution (3.19) for the density contrast. Note that we use an upper limit of infinity for the integration and confirm that it is not different from the integration result with an upper limit of $\delta=1$ from the requirement of avoiding separate Universes [Carr:1974nx, Yokoyama:1995ex]. A recent work [Kopp:2010sh] argued that this upper limit was a result of the geometry of the selected slicing rather than as a consequence of avoiding separate Universes, which instead led to a divergence of density fluctuations on the superhorizon scale. With the PBH mass given by Eq. (6.2), the fraction of PBHs at the formation time is
@@
where ${\rm erfc}(x)$ is the complementary error function (see appendix D for more details). The critical density contrast $\delta_{c}$ varies with the equation of state $\omega$. For radiation phase $\delta_{c}=0.453$. We determine $\delta_{c}$ with the $\delta_{c}-\omega$ relation revealed by the simulations in Ref. [Musco:2012au] (see section 7).
A more precise PBH mass formation was obtained from the numerical computations on the collapse of a relativistic fluid, which indicated a simple scaling law for the PBH mass distribution function near the critical threshold [Choptuik:1992jv, Evans:1994pj, Koike:1995jm]
@@
where the constant parameter $C=3.3$ and $b=0.36$ for the formation of PBH during radiation domination [Byrnes:2018clq]. Note that these parameters may be affected by the phase of cosmic matter. Then the PBH energy fraction is given by
@@
$s_{f}a_{f}^{3}=s_{0}a_{0}^{3},\penalty 10000\ \penalty 10000\ {\rm with}\penalty 10000\ \penalty 10000\ s(T)=\frac{2\pi^{2}}{45}g_{*,s}T^{3},$ (6.9)
where the current Universe entropy density $s_{0}=2891.2\penalty 10000\ \rm cm^{-3}$. Together with Eqs. (6.1)-(6.3) we obtain the the current PBH relic density abundance
@@
where the Hubble expansion rate is $H_{0}=100h\penalty 10000\ \mathrm{km\penalty 10000\ s}^{-1}\penalty 10000\ \mathrm{Mpc}^{-1}$ with $h=0.68$ from the Planck observation [Planck:2018vyg]. With the relation $M=\gamma_{p}M_{H}$ and the approximation $g_{*}(T)\simeq g_{*,s}(T)$, we reproduce $\Omega_{\rm PBH}$ given in Ref. [Carr:2009jm]. We observe that the PBH energy fraction at the formation should be of order $\beta(M)\sim 10^{-9}$ so that the PBH relic abundance today does not exceed the required dark matter relic abundance. Thus the variance of the density contrast should be small and the probability density distribution should be concentrated so that the probability for the production of large density contrasts is exponentially suppressed in the tail of the probability density distribution.
@@ 6.2 Density fluctuation from inflation-induced curvature perturbation
Let’s first briefly review the PBH production in the inflation scenario. For the density fluctuations from the curvature perturbations during inflation, the variance is in general estimated by Eq. (3.17). On comoving hypersurfaces, the density perturbation can be related to the curvature perturbation by [Green:2004wb]
@@
where the amplitude $A_{\mathcal{R}_{c}}$ is a constant that depends on the inflation physics, and $k_{*}$ is the reference scale. Current CMB and LSS observations on large scales ($k\lesssim{\rm Mpc^{-1}}$) have restricted the primordial curvature power spectrum at the order of around $10^{-9}$. To have a sizable PBH formation for LIGO observation, the primordial curvature power spectrum should be of order $\mathcal{P}_{\mathcal{R}_{c}}(k)\sim 10^{-2}-10^{-1}$ (corresponding to a variance in the range $10^{-4}\lesssim\sigma^{2}\lesssim 10^{-2}$ which gives $\beta\sim 10^{-9}-10^{-6}$, see Fig. 11 in appendix D), which requires a blue-tilted spectral index $n_{s}\simeq 1.85$ at scale $k\sim 10^{6}\penalty 10000\ \rm Mpc^{-1}$ (corresponding to a PBH mass $M\sim 30M_{\odot}$) [Sasaki:2018dmp, Byrnes:2018clq]. On the other hand, the CMB observations have constrained the spectral index to be red-tilted, $n_{s}=0.968\pm 0.006$ at the scale $k_{*}=0.05\penalty 10000\ \rm Mpc^{-1}$. The relatively large blue-tilted spectral index at a small scale cannot be realized in the single-field slow-roll inflation but may be achieved in the running mass inflation or multi-scalar inflation models, in which the scale-dependent spectral index can be realized (see Refs. [Stewart:1997wg, Drees:2011hb, Alabidi:2012ex, Silk:1986vc, Randall:1995dj, GarciaBellido:1996mdl, Clesse:2015wea] for example).